# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v2

@@ Abstract
Timely updating of Internet of Things (IoT) data is crucial for immersive vehicular metaverse services. However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder continuous collection of high-quality data. To address these issues, we propose an immersion-aware model trading framework that facilitates data provision for services while ensuring privacy through federated learning (FL). Specifically, we first develop a novel multi-dimensional metric, the immersion of model (IoM), which assesses model value comprehensively by considering freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training. Then, we design an incentive mechanism to incentivize metaverse users (MUs) to contribute high-value models under resource constraints. The trading interactions between MSPs and MUs are modeled as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains, where MSPs as leaders determine rewards, while MUs as followers optimize resource allocation. Furthermore, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process. To solve this, we develop a fully distributed dynamic reward algorithm based on deep reinforcement learning, without accessing any private information about MUs and other MSPs. Experimental results demonstrate that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
Index Terms:
Equilibrium problem with equilibrium constraints, immersion-aware, incentive mechanism, resource allocation, vehicular metaverse.
@@ I Introduction
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse [2, 3] has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies. Market report [4] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030. Automakers such as BMW have increasingly invested in AR technology. As shown in Fig. 2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you. Moreover, Nissan’s [5] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system with the intention of improving passengers’ parking routines and trip-end experiences.
@@
The ability to capture information from the “real” world, particularly the ability to collect and process massive data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse. Meanwhile, the data must be processed and presented in a meaningful, responsive, and appropriately protected way. Technically, a high-quality experience with AR services relies on accurate detection and classification of real-world objects (e.g., cars and pedestrians) under complex conditions [6]. To achieve this goal, sufficient valid data needs to be collected and processed in depth to detect and classify objects accurately. Therefore, it is essential to focus on effectively collecting, processing, and protecting the data that supports a safer and more enjoyable driving experience.
@@
Motivations. The widely used data collection method, as adopted by Nissan [7] and studied in [8], involves gathering massive amounts of data through vehicle sensors, cameras, and roadside devices, and all data being processed centrally. While this approach has been effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, the data synchronization will be hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [9]. Also, delays can severely impact the MU’s experience and cause dizziness [10]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs.
@@
Federated learning (FL) has been adopted in prior work [11, 12, 6, 13] to enable collaborative model training without sharing raw data (e.g., sensor/imaging data from vehicles). In this approach, learning models are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden. However, MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead. To overcome this, incentive mechanisms using strategies such as contract theory [14], Stackelberg game [15, 16] and multi-winner sealed-bid auction [17], have been proposed to encourage MUs to contribute models. However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the model for MSPs. Furthermore, most of these solutions are designed for a single MSP and do not consider the joint optimization of MUs’ limited computational and communication resources. As a result, an efficient and privacy-preserving framework for data synchronization is needed to improve the immersive AR experience in the vehicular metaverse.
@@
To address the above research gaps, we propose an immersion-aware model trading framework for a multi-MSP, multi-MU vehicular metaverse. This framework is designed to incentivize MUs to become active contributors by providing models tailored to the specific needs of MSPs. However, MUs have different sampling costs and limited computational and communication resources, while MSPs differ in their model preferences and compete with each other. These inherent asymmetries and competitive dynamics make it challenging to establish an efficient model trading ecosystem. This leads to two key research questions that underpin our work:
@@
  • How can MUs be effectively incentivized to contribute high-value learning models that benefit vehicular metaverse services for diverse MSPs?
@@
  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of learning models in the vehicular metaverse?
@@
Proposed Framework and Contributions. As depicted in Fig. 2, our model trading framework consists of four components: FL mechanism, metric design, game modeling, and dynamic algorithms. Specifically, we first design a new metric called the immersion of the learning model (IoM) to evaluate the value of learning models contributed by MUs to MSPs. This metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data. Building on this metric, we propose an immersion-aware incentive mechanism that aligns the interests of both MUs and MSPs. Then, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC), which is a hierarchical optimization problem with equilibria at two levels [18]. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. In summary, the key contributions of this work are as follows.
  • Incentive mechanism design for immersion-aware model trading. From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value learning models tailored to the specific demands of MSPs. To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU, multi-MSP vehicular metaverse environments.
  • Novel design of multi-dimensional metric. To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the learning model integrating four critical dimensions: the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data. Freshness is captured through age of information (AoI), while potential value is evaluated by the difference between model predictions and true labels. This metric enables fine-grained evaluation of model value under resource constraints and provides a basis for incentive and decision-making strategies in model trading.
@@
  • Performance evaluation. We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Numerical results show that our proposed mechanism improves the IoM by $38.3\%$ and $37.2\%$, and reduces the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively, compared with state-of-the-art benchmarks.
The rest of this paper is organized as follows. Section II discusses the related work. In Section III, we present the system overview and design the immersion metric of the learning model. Section IV gives the game formulation, and Section V analyzes the existence of the equilibria at two levels. In Section VI, we give the detailed design of MDDR. Section VII shows numerical experiments to evaluate the framework performance, and finally Section VIII concludes the paper.
@@ II Related Work
In this section, we survey work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
@@ II-A Edge-Enabled Vehicular Metaverse Services
Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences such as AR navigation and real-time digital twins. To meet these demands, researchers have extensively explored edge-based strategies for task offloading and resource management. For instance, Feng et al. [19] proposed a resource allocation framework for AR-powered vehicular edge metaverses that aims to maximize overall system utility. Similarly, Khan et al. [20] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce transmission energy consumption and latency in resource-constrained environments. Beyond static offloading, Chen et al. [21] developed a multi-agent deep reinforcement learning approach for dynamic avatar task migration. In terms of security and privacy, Kang et al. [22] presented a cross-metaverse empowered dual pseudonym management framework to mitigate privacy leakage risks during the dynamic communications among vehicular edge metaverses. Moreover, an optimization framework for AR-assisted driving was proposed in [23] to enhance the robustness of metaverse maps against adversarial attacks. This framework maximizes the mean average precision (mAP) of adversarial patch detection while minimizing AR scene uplink latency and worst-case battery consumption. Recent studies have also explored applying FL to vehicular metaverse services as a means of preserving user privacy during distributed model training. These efforts, particularly in AR-enhanced metaverse environments, are discussed in the next section. To ensure service continuity and prevent user disengagement during unexpected failures, Qiu et al. [24] proposed deploying redundant backup virtual vehicle services and keeping the age of backup information to improve the reliability of services.
These works, from diverse perspectives, have significantly advanced the security, reliability, and performance of vehicular metaverse services, providing valuable insights for future research. Notably, distinct from existing studies, our work focuses on designing an incentive mechanism that transforms MUs into active contributors. These contributors efficiently provide high-value models to MSPs, thereby enhancing the overall service experience.
@@ II-B Federated Learning for AR in the Metaverse
Privacy preservation is a critical concern in vehicular metaverse services, particularly when training models using sensitive user data. Federated learning (FL) has emerged as a promising technique to enable collaborative model training without sharing raw data. For example, Chen et al. [11] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. However, they did not address how communication and computational resources should be allocated within the FL framework. To overcome this limitation, Zhou et al. [12] incorporated FL into mobile AR systems in the metaverse to enable collaborative model training. Their approach considers the trade-offs among energy consumption, execution latency, and model accuracy to meet varying demands across different application scenarios. To further improve learning efficiency, an adaptive and resource-efficient FL algorithm tailored for AR applications was proposed in [6]. This approach mitigates the effects of non-IID data distributions, reduces resource usage, and enhances the quality of experience. Additionally, Hazarika et al. [13] explored the integration of quantum computing with FL to provide a cost-efficient and adaptive solution for the dynamic nature of vehicular environments.
@@
The quality of metaverse services depends heavily on massive sensing data from the real world. To facilitate this, incentive mechanisms are widely employed to motivate MUs to contribute data and participate in model construction, thereby enhancing their service experience. For instance, Zhang et al. [25] proposed a vehicle-assisted data sensing framework that incentivizes vehicles to contribute sensing data to metaverse service, improving the driving experience for vehicle users. Similarly, Lin et al. [26] introduced a contract-based incentive design aimed at optimizing the long-term profits of digital twin service providers, addressing service congestion caused by stochastic demand responses. Xu et al. [27] developed a metaverse-based unmanned aerial vehicle (UAV) swarm system that uses digital twins and semantic communication to enhance data synchronization and reduce communication latency. Moreover, Lotfi et al. [28] introduced VMGuard, a novel four-layer security framework for the vehicular metaverse that protects against data poisoning attacks by using a reputation-based incentive mechanism. The framework leveraged user feedback and subjective logic modeling to assess the trustworthiness of devices. However, while these works successfully incentivize MUs to contribute raw data, they also pose a risk to user privacy due to potential data leakage.
@@
To address the privacy concern, Kang et al. [14] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory to incentivize data sharing. They utilized a hierarchical cross-chain architecture with a main chain and multiple subchains to perform decentralized, privacy-preserving, and secure data training in both virtual and physical spaces. Baccour et al. [15] proposed a dual game-theoretic framework for federated meta-learning (FML) in metaverse services. They introduced a reputation system and a Stackelberg game-based incentive mechanism to enhance data privacy, minimize energy costs, and improve FML efficacy, outperforming traditional clustering methods in training performance and metaverse utility. Li et al. [16] designed a satisfaction metric that considers data size, AoI, and training latency. They integrated this metric into utility functions to incentivize node participation in FL while optimizing resource allocation. However, these works are designed for a single MSP scenario, limiting scalability and applicability in multi-server scenarios commonly encountered in large-scale metaverse environments.
Building on the concept of multiple MSPs, Zhang et al. [17] proposed a privacy-preserving auction mechanism for learning model as an NFT in Blockchain-driven Metaverse. They employed an imperfect information Stackelberg game to optimize the strategies of MUs and MSPs, and introduced a multi-winner sealed-bid auction mechanism to enhance the quality of FL-NFT. However, this approach primarily focuses on modeling NFTs and overlooks the joint optimization of computational and communication resources. Moreover, it fails to evaluate the model’s value from multiple dimensions.
Inspired by the above existing works, this study focuses on multi-MSP and multi-MU scenarios in vehicular metaverse services. By assigning MUs the dual roles of service consumers and model contributors, we build a low-cost, privacy-preserving data support framework. This allows MUs to enjoy immersive in-vehicle services (e.g., AR navigation, digital twin interactions) while participating as distributed model contributors in service optimization. This framework continuously incentivizes MUs to contribute high-value learning models for metaverse services, thereby enhancing the overall service experience.
@@
Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..

### v3

@@ Abstract
Timely updating of Internet of Things (IoT) data is crucial for achieving immersion in vehicular metaverse services. However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder the continuous collection of high-quality data. To address these challenges, we propose an immersion-aware model trading framework that enables efficient and privacy-preserving data provisioning through federated learning (FL). Specifically, we first develop a novel multi-dimensional evaluation metric for the immersion of models (IoM). The metric considers i) the freshness and accuracy of the local model, and ii) the amount and potential value of raw training data. Building on the IoM, we design an incentive mechanism to encourage metaverse users (MUs) to participate in FL by providing local updates to MSPs under resource constraints. The trading interactions between MSPs and MUs are modeled as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains, where MSPs as leaders determine rewards, while MUs as followers optimize resource allocation. To ensure privacy and adapt to dynamic network conditions, we develop a distributed dynamic reward algorithm based on deep reinforcement learning, without acquiring any private information from MUs and other MSPs. Experimental results show that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively. These findings validate the effectiveness of our approach in incentivizing MUs to contribute high-value local models to MSPs, providing a flexible and adaptive scheme for data provisioning in vehicular metaverse services.
@@ I Introduction
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies. Market report [2] projected that the global automotive metaverse market would grow from \$$1.9$ billion in 2022 to \$$6.5$ billion by 2030. Automakers such as BMW are actively investing in AR technologies. As shown in Fig. 2, augmented information improves driving safety by displaying potential hazards obscured by the vehicle ahead. Moreover, Nissan’s [3] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system to improve passengers’ parking routines and trip-end experiences.
@@
The ability to capture information from the “real” world, particularly the ability to collect and process massive amounts of data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse. Meanwhile, the data must be processed and presented in a meaningful and responsive manner with adequate privacy protection. Technically, a high-quality AR experience relies on accurately detecting and classifying real-world objects (e.g., cars and pedestrians) under complex conditions [4]. To achieve this goal, sufficient valid data need to be collected and deeply processed to detect and classify objects accurately. Therefore, it is essential to focus on effectively collecting, processing, and protecting the data that supports a safer and more enjoyable driving experience.
@@
Motivations. The widely used data collection method, as adopted by Nissan [5] and studied in [6], involves gathering massive amounts of data through vehicle sensors, cameras, and roadside devices, and all data being processed centrally. While this approach is effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, data synchronization is hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [7]. Also, delays can severely impact the MU’s experience and cause dizziness [8]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs.
@@
Federated learning (FL) has been adopted in prior work [9, 10, 4, 11] to enable collaborative model training without sharing raw data (e.g., sensor/imaging data from vehicles). In this approach, local updates are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden. However, MUs are typically self-interested and are reluctant to share local models with MSPs due to the additional computation, communication, and energy overhead. To overcome this issue, incentive mechanisms using strategies such as contract theory [12], Stackelberg game [13, 14] and multi-winner sealed-bid auction [15], have been proposed to encourage MUs to contribute local models. However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the models for MSPs. Furthermore, most of these solutions are designed for a single MSP and do not consider the joint optimization of MUs’ limited computational and communication resources. As a result, an efficient and privacy-preserving framework for data synchronization is needed to improve the immersive AR experience in the vehicular metaverse.
@@
To address the above research gaps, we propose an immersion-aware model trading framework for a multi-MSP, multi-MU vehicular metaverse. This framework is designed to incentivize MUs to become active contributors by providing local models tailored to the specific needs of MSPs. However, MUs have different sampling costs and limited computational and communication resources, while MSPs differ in their model preferences and compete with each other. These inherent asymmetries and competitive dynamics make it challenging to establish an efficient model trading ecosystem. This leads to two key research questions that underpin our work:
@@
  • How can MUs be effectively incentivized to contribute high-value local models that benefit vehicular metaverse services for diverse MSPs?
@@
  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of local models in the vehicular metaverse?
@@
Proposed Framework and Contributions. As depicted in Fig. 2, our model trading framework consists of four components: FL mechanism, metric design, game modeling, and dynamic algorithms. Specifically, this work first designs a new metric called the immersion of the learning model (IoM) to evaluate the value of local models contributed by MUs to MSPs. This metric jointly considers the freshness and accuracy of the local model, as well as the amount and potential value of raw training data. Building on this metric, we propose an immersion-aware incentive mechanism that aligns the interests of both MUs and MSPs. Then, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC), which is a hierarchical optimization problem with equilibria at two levels [16]. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. In summary, the key contributions of this work are as follows.
  • Incentive mechanism design for immersion-aware model trading. From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value local models tailored to the specific demands of MSPs. To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU and multi-MSP vehicular metaverse environments.
  • Novel design of multi-dimensional metric. To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the local model integrating four critical dimensions: the freshness and accuracy of the local model, as well as the amount and potential value of raw training data. Freshness is captured through age of information (AoI), while potential value is evaluated by the difference between model predictions and true labels. This metric enables fine-grained evaluation of model value under resource constraints and provides a basis for incentive and decision-making strategies in model trading.
@@
  • Performance evaluation. Extensive numerical simulations are conducted based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Numerical results show that our proposed mechanism improves the IoM by $38.3\%$ and $37.2\%$, and reduces the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively, compared with state-of-the-art benchmarks.
The rest of this paper is organized as follows. Section II discusses the related work. In Section III, we present the system overview and design the immersion metric of the local model. Section IV gives the game formulation, and Section V analyzes the existence of the equilibria at two levels. In Section VI, we give the detailed design of MDDR. Section VII shows numerical experiments to evaluate the framework performance, and finally, Section VIII concludes the paper.
@@ II Related Work
In this section, we discuss the work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
@@ II-A Edge-Enabled Vehicular Metaverse Services
Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences, such as AR navigation and real-time digital twins. To meet these demands, researchers have extensively explored edge-based strategies for task offloading and resource management. For instance, Feng et al. [17] proposed an innovative resource allocation framework for AR-empowered vehicular edge metaverse, which significantly improves data utility and reduces vehicle energy consumption through multi-dimensional optimization. Similarly, Khan et al. [18] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce both transmission energy consumption and latency in resource-constrained environments. To cope with dynamic network conditions, Tong et al. [19] proposed an attribute-aware auction mechanism for resource allocation, and applied a GPT-based deep reinforcement learning algorithm to adjust the auction clocks. Regarding security and privacy, Kang et al. [20] presented a cross-metaverse empowered dual pseudonym management framework to mitigate privacy leakage risks during the dynamic communications among vehicular edge metaverses. Moreover, a novel four-layer security framework was designed to protect vehicular metaverse systems from data poisoning attacks [21].
In addition to theoretical models and optimization frameworks for integrating autonomous vehicles into the metaverse, several studies have proposed practical decision-making approaches to evaluate and select among alternative implementation strategies. For example, Deveci et al. [22] proposed a framework that combines a fuzzy logarithmic weighting method with a similarity-based ranking technique to assess four traffic safety solutions enabled by the metaverse. Their results show that public transportation is the most practical and scalable option. Moreover, they applied a decision-making method based on Q-rung orthopair fuzzy sets to rank personal mobility alternatives, offering practical guidance for system planning in [23]. Gokasar et al. [6] introduced a two-stage fuzzy decision-making framework to evaluate integration options for self-powered sensors in autonomous vehicles. The study concludes that using these sensors for real-time traffic management in the metaverse is the most effective solution.
These works, from diverse perspectives, have significantly advanced the security, reliability, and performance of vehicular metaverse services, providing valuable insights for future research. Notably, distinct from existing studies, our work focuses on designing an incentive mechanism that transforms MUs into active contributors. These contributors efficiently provide high-value local models to MSPs, thereby enhancing the overall service experience.
@@ II-B Federated Learning for AR in the Metaverse
Privacy preservation is critical in vehicular metaverse services, particularly when training models using sensitive user data. Federated learning (FL) has emerged as a promising technique to enable collaborative model training without sharing raw data. For example, Chen et al. [9] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. However, they did not address how communication and computational resources should be allocated within the FL framework.
To overcome this limitation, Zhou et al. [10] developed an FL-assisted MAR system via non-orthogonal multiple access for the metaverse. Furthermore, a resource allocation algorithm was designed to balance accuracy and energy consumption. To further improve learning efficiency, an adaptive and resource-efficient FL algorithm tailored for AR applications was proposed in [4]. This approach mitigates the effects of non-IID data distributions, reduces resource usage, and enhances the quality of experience. Additionally, Hazarika et al. [11] explored the integration of quantum computing with FL to provide a cost-efficient and adaptive solution for the dynamic nature of vehicular environments. Nadimi et al. [24] proposed a multi-modal, multi-task federated foundation model (FedFMs) architecture. It combines the privacy-preserving benefits of FL with the representation capabilities of foundation models to address sensor diversity, personalized interactions, and device heterogeneity in extended reality systems.
@@
TABLE I: Incentive Mechanisms for Data Synchronization
Ref.
MSPs
MUs
Theory
Method
Main
Metric
Data
Privacy
[12] $\times$
Contract theory
under PT
AoI
[13] $\times$
Dual
game-theoretic
Reputation
scores
[14] $\times$ Stackelberg game
Model quality
and latency
[15] IISG and PMS-AM
Model quality
and freshness
[25]
Coalition formation,
Stackelberg game
Volume of
sensing data
$\times$
[26]
Deep learning
based auction
Quality,
reliability, latency
$\times$
[27] Contract theory Computing delay $\times$
[28] Stackelberg game
Energy
consumption
$\times$
[29]
Multi-objective
optimization
Peak signal-
to-noise ratio
Our
work
EPEC IoM
@@
Immersive metaverse services depend on high-quality real-world sensing data. However, the process of collecting such data is often costly and may raise privacy concerns. To this end, various incentive mechanisms have been proposed to encourage MUs to contribute data and participate in model training. For instance, Zhang et al. [25] introduced a vehicle-assisted data sensing framework that incentivizes vehicles to upload sensing data, enhancing the driving experience. Xu et al. [26] developed a UAV swarm system using digital twins and semantic communication to improve data synchronization and reduce latency. Lin et al. [27] designed an incentive-based congestion control scheme for digital twin edge networks, which addresses the stochastic service demand and the long-term provider profit by combining Lyapunov optimization and contract theory. Zhao et al. [28] proposed a physical reality incentive mechanism based on a two-stage Stackelberg game to motivate MUs to collect physical reality sensing data for the metaverse. Although these studies successfully encourage raw data contributions, they also pose a risk to user privacy due to potential data leakage.
@@
To address privacy concerns, Kang et al. [12] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory (PT) to incentivize data sharing. They utilized a hierarchical cross-chain architecture with a main chain and multiple subchains to perform decentralized, privacy-preserving, and secure data training in both virtual and physical spaces. Baccour et al. [13] designed a dual game-theoretic framework for federated meta-learning (FML), incorporating a reputation system and Stackelberg game to enhance privacy, reduce energy consumption, and improve model performance. Li et al. [14] designed a new metric named “Satisfaction” to balance training latency and model quality. They integrated this metric into utility functions to incentivize node participation in FL while optimizing resource allocation. However, these works are designed for a single MSP scenario, limiting scalability and applicability in multi-server scenarios commonly encountered in large-scale metaverse environments.
To fill this gap, Zhang et al. [15] employed an imperfect information Stackelberg game (IISG) to optimize the strategies of MUs and MSPs, and introduced a privacy-preserving multi-winner sealed-bid auction mechanism (PMS-AM) to enhance the quality of non-fungible token (NFT) with FL assistance. Esmail et al. [29] proposed a metaverse FL and semantic communication (MFSC) framework that matches IoT device capabilities to training tasks, optimizing digital twin quality through adaptive resource allocation. Nonetheless, existing incentive criteria are often simplistic and fail to evaluate the values of contributed models comprehensively.
Unlike these prior works, our study focuses on vehicular metaverse services in multi-MSP and multi-MU scenarios. We propose a novel metric, IoM, which evaluates model value from multiple dimensions. Based on this metric, we design an incentive mechanism that better motivates MUs to contribute high-value local models, thereby improving the quality and user experience of metaverse services. A detailed comparison of these approaches is provided in Table I, highlighting the key contributions of this study.
@@
Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..